# Deep Learning, Machine Learning - Digital Signal and Image Processing: From Theory to Application

### v1

@@ 3.2 Relationship Between DFT and DTFT
The Discrete Fourier Transform (DFT) can be seen as a sampled version of the Discrete-Time Fourier Transform (DTFT) [136, 20]. While the DTFT applies to an infinite-length sequence and produces a continuous frequency spectrum, the DFT is used for finite-length signals and gives a discrete set of frequency samples.
In practical applications, we often work with finite signals (e.g., audio signals, sensor data, etc.), and thus, the DFT is more commonly used. The DFT samples the frequency spectrum at $N$ equally spaced points between 00 and $2\pi$, where $N$ is the length of the signal [20, 155].
@@ 3.3.1 Periodicity
Example: Periodicity of the DFT Let’s take a sequence $x[n]=\{1,0,0,0\}$ of length 4 and compute its DFT. The result will be periodic with a period of 4, meaning that $X[k]$ repeats after every 4 frequency components [127].
@@ 3.6 Fast Fourier Transform (FFT) Algorithm
In this section, we will discuss one of the most powerful algorithms in Digital Signal Processing (DSP), the Fast Fourier Transform (FFT) [23, 136]. The FFT is an efficient method for computing the Discrete Fourier Transform (DFT) and its inverse. The FFT dramatically reduces the time complexity, making it possible to perform spectral analysis, filtering, and compression on large datasets in real-time applications [155].
@@ 3.6.1 Introduction to FFT
The Discrete Fourier Transform (DFT) is widely used for converting signals from the time domain to the frequency domain [186]. However, the direct computation of the DFT requires $O(N^{2})$ operations for a signal of length $N$, which is computationally expensive for large datasets.
The Fast Fourier Transform (FFT) algorithm solves this problem by reducing the computational complexity to $O(N\log N)$ [137]. This improvement allows for the efficient analysis of large signals and real-time applications such as audio processing, communications, and image compression [20].
@@ 3.6.2 Cooley-Tukey Algorithm
Specifically, the Cooley-Tukey algorithm works by factorizing the DFT of a composite number $N$ (where $N$ is not a prime number) into multiple smaller DFTs. This significantly decreases the computational burden, as smaller DFTs can be computed more efficiently [156].
@@ 3.6.3 Radix-2 FFT Algorithm
The Radix-2 FFT algorithm is a specific implementation of the Cooley-Tukey algorithm when $N$, the length of the input signal, is a power of 2 [23]. This is the most common case in practice because it allows for highly efficient computation of the FFT [136].
The Radix-2 FFT works by recursively dividing the input sequence into two parts: one consisting of the even-indexed elements and the other consisting of the odd-indexed elements. The DFTs of these two sequences are then combined using symmetry properties of the DFT [156].
@@ 3.6.4 Applications of FFT
The FFT is one of the most widely used algorithms in various fields, including signal processing, communications, and multimedia [186]. Here, we will discuss a few key applications where the FFT plays a critical role.
@@
Spectral analysis involves examining the frequency content of a signal to detect dominant frequencies and patterns. In audio processing, for example, the FFT is used to analyze sound waves and detect pitch, harmonics, and noise [137]. In radar and sonar, the FFT helps detect objects by analyzing the frequency content of reflected signals [163].
@@
Filtering is one of the core applications of the FFT [155]. Instead of performing convolution in the time domain, which can be computationally expensive, signals are transformed into the frequency domain using the FFT, filtered by multiplication with a frequency-domain filter, and then transformed back to the time domain using the inverse FFT (IFFT) [58].
@@
FFT is also an essential tool in signal and image compression techniques [37]. For example, in JPEG compression, the FFT (or more commonly, the Discrete Cosine Transform, which is closely related) is used to transform the image into the frequency domain. In the frequency domain, high-frequency components, which often correspond to noise or insignificant details, are discarded or quantized, reducing the data size without significantly affecting image quality [127].
@@ 4.1 Definition of Z-Transform
The Z-transform is one of the most important tools in digital signal processing (DSP) and control systems [136]. It extends the Fourier transform for analyzing discrete-time signals and systems, providing insights into the behavior of systems in both the time and frequency domains. The Z-transform converts a discrete-time signal $x[n]$ into a complex function of a variable $z$, defined as:
@@
  • $z$ is a complex variable, often written as $z=re^{j\omega}$, where $r$ is the magnitude and $\omega$ is the angular frequency [186].
The Z-transform is extremely versatile and generalizes the concept of the Discrete-Time Fourier Transform (DTFT) [137]. While the DTFT is limited to the unit circle (i.e., $z=e^{j\omega}$), the Z-transform provides a broader perspective by considering all possible values of $z$.
@@ 4.2 Region of Convergence (ROC)
The Region of Convergence (ROC) is a critical concept in the analysis of the Z-transform [52]. The ROC defines the set of values of the complex variable $z$ for which the Z-transform sum converges. Understanding the ROC is essential because it determines whether a given Z-transform corresponds to a valid time-domain signal and helps to assess important system properties like stability and causality [127].
@@ 4.2.1 Why ROC is Important
  • Two-Sided Signals: For signals that are neither causal nor anti-causal, the ROC is a ring-shaped region between poles [137].
@@ 4.3 Inverse Z-Transform
The Inverse Z-transform is used to recover the original discrete-time signal $x[n]$ from its Z-domain representation $X(z)$ [136, 155]. While the Z-transform maps time-domain sequences into the $z$-domain, the inverse process allows us to return from the $z$-domain back to the time domain [186].
@@ 4.3.1 Methods for Inverse Z-Transform
There are several methods for finding the inverse Z-transform, depending on the nature of the function $X(z)$ [58]:
  • Partial Fraction Expansion: This method is commonly used when $X(z)$ is a rational function (i.e., a ratio of polynomials). By breaking $X(z)$ into simpler components, we can use Z-transform tables to find the corresponding time-domain signal [127].
@@
  • Using Z-Transform Tables: In many cases, pre-computed Z-transform tables provide a straightforward way to find the inverse Z-transform by matching the form of $X(z)$ to known results [136].
@@ 4.4 Python Example: Computing the Z-Transform
While Python does not have a built-in function for computing the Z-transform directly, we can implement it manually or use symbolic computation libraries like sympy [123]. In the example below, we use sympy to compute the Z-transform and inverse Z-transform of a simple sequence.
@@ 4.5 System Analysis Using the Z-Transform
The Z-transform is also used to analyze discrete-time systems [155]. By converting a system’s difference equation into the Z-domain, we can easily solve for its behavior and analyze properties like stability, frequency response, and transient response [136].
@@ 4.5.2 Stability and the ROC
For a system to be stable, all poles of its transfer function must lie within the unit circle (i.e., $|z|<1$) [52]. The ROC must include the unit circle for the system to be stable.
@@ 4.7 System Function and System Analysis
In Digital Signal Processing (DSP), the system function provides a complete description of a linear time-invariant (LTI) system in the Z-domain [136]. The system function, denoted as $H(z)$, is the Z-transform of the system’s impulse response, $h[n]$ [155]. By analyzing the system function, we can gain critical insights into the behavior of the system, including stability, frequency response, and phase response [186].
@@ 4.7.2 Poles and Zeros of the System Function
The poles and zeros of the system function $H(z)$ are essential for understanding the behavior of the system [137]. Zeros are the values of $z$ that make $H(z)=0$, and poles are the values of $z$ that make $H(z)$ undefined (where the denominator $D(z)=0$).
@@ 4.7.3 Stability of a System
The stability of an LTI system depends on the location of the poles of the system function $H(z)$ in the Z-plane. For a system to be stable, all poles must lie inside the unit circle ($|z|<1$) in the Z-plane [52]. If any pole lies outside or on the unit circle, the system is unstable.
@@ 4.8 Applications of Z-Transform
The Z-transform is a powerful mathematical tool used extensively in the analysis and design of digital systems [58]. It provides insights into the behavior of discrete-time systems in both the time and frequency domains. Here are a few key applications of the Z-transform in DSP [155].
@@ 4.8.1 Digital Filter Design
One of the primary applications of the Z-transform is in the design of digital filters [136]. Filters are used to selectively enhance or suppress specific frequency components in a signal. By analyzing the Z-transform of the filter’s impulse response, we can design and analyze filters such as low-pass, high-pass, band-pass, and band-stop filters.
@@ 4.8.2 Control Systems Analysis
The Z-transform is widely used in the analysis of discrete-time control systems [127]. In control theory, the Z-transform helps analyze system stability, transient response, and steady-state behavior. By examining the system’s poles in the Z-plane, control engineers can assess whether the system will reach a stable equilibrium and how it will react to changes in input.
@@ 4.8.3 Time-Domain and Frequency-Domain Analysis
The Z-transform provides a bridge between the time domain and the frequency domain [136]. By analyzing the Z-transform of a signal or system, engineers can understand how a system behaves over time and across different frequencies [155]. This is particularly useful in applications like audio processing, communications, and image compression, where signals are analyzed and processed in both domains [186].
@@ 5.1 Types of Digital Filters
Digital filters are essential components in digital signal processing, used to manipulate or extract information from discrete-time signals [136]. They are algorithms or devices that operate on a digital input signal to produce a digital output signal. Digital filters are broadly categorized into two main types: Finite Impulse Response (FIR) filters and Infinite Impulse Response (IIR) filters. Each type has unique characteristics, advantages, and applications.
@@ 5.1.1 Finite Impulse Response (FIR) Filters
FIR filters have an impulse response that is of finite duration, meaning it settles to zero in finite time. The general form of an FIR filter is given by the convolution sum [155]:
@@
  • $N$ is the number of taps (filter order + 1).
@@
  • Always Stable: Since FIR filters do not have feedback elements, they are inherently stable [136].
  • Linear Phase Response: FIR filters can be designed to have a linear phase response, which preserves the wave shape of signals within the passband [184].
  • Ease of Design: FIR filters are generally easier to design using standard methods [184].
@@ 5.1.2 Infinite Impulse Response (IIR) Filters
IIR filters have an impulse response that theoretically lasts forever, as they use feedback from output to input [136]. The general form of an IIR filter is given by the difference equation:
@@
  • Potentially Unstable: Due to feedback, stability must be carefully considered [155].
  • Efficient: IIR filters can achieve a desired response with a lower filter order than FIR filters [184].
  • Non-linear Phase Response: IIR filters generally do not have linear phase characteristics [136].
@@ 5.2 FIR Filter Design Techniques
Designing FIR filters involves determining the filter coefficients $b_{k}$ that meet specific frequency response characteristics. Two commonly used techniques for FIR filter design are the Window Method and the Frequency Sampling Method [155].
@@ 5.2.1 Window Method
The Window Method is a straightforward and widely used approach for designing FIR filters. The main idea is to truncate the infinite impulse response of an ideal filter to a finite length by applying a window function [184]. This process introduces ripples in the frequency response due to the Gibbs phenomenon, but by choosing appropriate window functions, we can control the trade-off between main-lobe width and side-lobe levels [155].
@@
The general steps for designing an FIR filter using the window method are as follows [184]:
@@
    Choose a Window Function: Select an appropriate window function $w[n]$ (e.g., Rectangular, Hamming, Hanning, Blackman, or Kaiser window) that meets the design specifications [155].
@@
Different window functions affect the frequency response differently [184]. Here are some commonly used windows:
@@
    Simplest window with poor frequency response (high side-lobes) [184].
@@
    Provides better side-lobe attenuation than the Rectangular window [184].
@@
    Similar to Hamming but with different side-lobe characteristics [184].
@@
    Offers better side-lobe attenuation at the expense of a wider main lobe [184].
@@
    Where $I_{0}$ is the zero-th order modified Bessel function, and $\beta$ is a parameter that controls the trade-off between main-lobe width and side-lobe level [155].
@@
We can use Python to compute the filter coefficients and visualize the frequency response [203].

### v2

@@ 3.2 Relationship Between DFT and DTFT
The Discrete Fourier Transform (DFT) can be seen as a sampled version of the Discrete-Time Fourier Transform (DTFT) [oppenheim2010dsp, bracewell2000fourier]. While the DTFT applies to an infinite-length sequence and produces a continuous frequency spectrum, the DFT is used for finite-length signals and gives a discrete set of frequency samples.
In practical applications, we often work with finite signals (e.g., audio signals, sensor data, etc.), and thus, the DFT is more commonly used. The DFT samples the frequency spectrum at $N$ equally spaced points between 0 and $2\pi$, where $N$ is the length of the signal [bracewell2000fourier, proakis1996digital].
@@ 3.3.1 Periodicity
Example: Periodicity of the DFT Let’s take a sequence $x[n]=\{1,0,0,0\}$ of length 4 and compute its DFT. The result will be periodic with a period of 4, meaning that $X[k]$ repeats after every 4 frequency components [mitra2002digital].
@@ 3.6 Fast Fourier Transform (FFT) Algorithm
In this section, we will discuss one of the most powerful algorithms in Digital Signal Processing (DSP), the Fast Fourier Transform (FFT) [brigham1988fast, oppenheim2010dsp]. The FFT is an efficient method for computing the Discrete Fourier Transform (DFT) and its inverse. The FFT dramatically reduces the time complexity, making it possible to perform spectral analysis, filtering, and compression on large datasets in real-time applications [proakis1996digital].
@@ 3.6.1 Introduction to FFT
The Discrete Fourier Transform (DFT) is widely used for converting signals from the time domain to the frequency domain [smith1997scientist]. However, the direct computation of the DFT requires $O(N^{2})$ operations for a signal of length $N$, which is computationally expensive for large datasets.
The Fast Fourier Transform (FFT) algorithm solves this problem by reducing the computational complexity to $O(N\log N)$ [oppenheim1997signals]. This improvement allows for the efficient analysis of large signals and real-time applications such as audio processing, communications, and image compression [bracewell2000fourier].
@@ 3.6.2 Cooley-Tukey Algorithm
Specifically, the Cooley-Tukey algorithm works by factorizing the DFT of a composite number $N$ (where $N$ is not a prime number) into multiple smaller DFTs. This significantly decreases the computational burden, as smaller DFTs can be computed more efficiently [rader1968discrete].
@@ 3.6.3 Radix-2 FFT Algorithm
The Radix-2 FFT algorithm is a specific implementation of the Cooley-Tukey algorithm when $N$, the length of the input signal, is a power of 2 [brigham1988fast]. This is the most common case in practice because it allows for highly efficient computation of the FFT [oppenheim2010dsp].
The Radix-2 FFT works by recursively dividing the input sequence into two parts: one consisting of the even-indexed elements and the other consisting of the odd-indexed elements. The DFTs of these two sequences are then combined using symmetry properties of the DFT [rader1968discrete].
@@ 3.6.4 Applications of FFT
The FFT is one of the most widely used algorithms in various fields, including signal processing, communications, and multimedia [smith1997scientist]. Here, we will discuss a few key applications where the FFT plays a critical role.
@@
Spectral analysis involves examining the frequency content of a signal to detect dominant frequencies and patterns. In audio processing, for example, the FFT is used to analyze sound waves and detect pitch, harmonics, and noise [oppenheim1997signals]. In radar and sonar, the FFT helps detect objects by analyzing the frequency content of reflected signals [richards2014fundamentals].
@@
Filtering is one of the core applications of the FFT [proakis1996digital]. Instead of performing convolution in the time domain, which can be computationally expensive, signals are transformed into the frequency domain using the FFT, filtered by multiplication with a frequency-domain filter, and then transformed back to the time domain using the inverse FFT (IFFT) [gonzalez2002digital].
@@
FFT is also an essential tool in signal and image compression techniques [cohen2019biomedical]. For example, in JPEG compression, the FFT (or more commonly, the Discrete Cosine Transform, which is closely related) is used to transform the image into the frequency domain. In the frequency domain, high-frequency components, which often correspond to noise or insignificant details, are discarded or quantized, reducing the data size without significantly affecting image quality [mitra2002digital].
@@ 4.1 Definition of Z-Transform
The Z-transform is one of the most important tools in digital signal processing (DSP) and control systems [oppenheim2010dsp]. It extends the Fourier transform for analyzing discrete-time signals and systems, providing insights into the behavior of systems in both the time and frequency domains. The Z-transform converts a discrete-time signal $x[n]$ into a complex function of a variable $z$, defined as:
@@
  • $z$ is a complex variable, often written as $z=re^{j\omega}$, where $r$ is the magnitude and $\omega$ is the angular frequency [smith1997scientist].
The Z-transform is extremely versatile and generalizes the concept of the Discrete-Time Fourier Transform (DTFT) [oppenheim1997signals]. While the DTFT is limited to the unit circle (i.e., $z=e^{j\omega}$), the Z-transform provides a broader perspective by considering all possible values of $z$.
@@ 4.2 Region of Convergence (ROC)
The Region of Convergence (ROC) is a critical concept in the analysis of the Z-transform [forouzan2016region]. The ROC defines the set of values of the complex variable $z$ for which the Z-transform sum converges. Understanding the ROC is essential because it determines whether a given Z-transform corresponds to a valid time-domain signal and helps to assess important system properties like stability and causality [mitra2002digital].
@@ 4.2.1 Why ROC is Important
  • Two-Sided Signals: For signals that are neither causal nor anti-causal, the ROC is a ring-shaped region between poles [oppenheim1997signals].
@@ 4.3 Inverse Z-Transform
The Inverse Z-transform is used to recover the original discrete-time signal $x[n]$ from its Z-domain representation $X(z)$ [oppenheim2010dsp, proakis1996digital]. While the Z-transform maps time-domain sequences into the $z$-domain, the inverse process allows us to return from the $z$-domain back to the time domain [smith1997scientist].
@@ 4.3.1 Methods for Inverse Z-Transform
There are several methods for finding the inverse Z-transform, depending on the nature of the function $X(z)$ [gonzalez2002digital]:
  • Partial Fraction Expansion: This method is commonly used when $X(z)$ is a rational function (i.e., a ratio of polynomials). By breaking $X(z)$ into simpler components, we can use Z-transform tables to find the corresponding time-domain signal [mitra2002digital].
@@
  • Using Z-Transform Tables: In many cases, pre-computed Z-transform tables provide a straightforward way to find the inverse Z-transform by matching the form of $X(z)$ to known results [oppenheim2010dsp].
@@ 4.4 Python Example: Computing the Z-Transform
While Python does not have a built-in function for computing the Z-transform directly, we can implement it manually or use symbolic computation libraries like sympy [meurer2017sympy]. In the example below, we use sympy to compute the Z-transform and inverse Z-transform of a simple sequence.
@@ 4.5 System Analysis Using the Z-Transform
The Z-transform is also used to analyze discrete-time systems [proakis1996digital]. By converting a system’s difference equation into the Z-domain, we can easily solve for its behavior and analyze properties like stability, frequency response, and transient response [oppenheim2010dsp].
@@ 4.5.2 Stability and the ROC
For a system to be stable, all poles of its transfer function must lie within the unit circle (i.e., $|z|<1$) [forouzan2016region]. The ROC must include the unit circle for the system to be stable.
@@ 4.7 System Function and System Analysis
In Digital Signal Processing (DSP), the system function provides a complete description of a linear time-invariant (LTI) system in the Z-domain [oppenheim2010dsp]. The system function, denoted as $H(z)$, is the Z-transform of the system’s impulse response, $h[n]$ [proakis1996digital]. By analyzing the system function, we can gain critical insights into the behavior of the system, including stability, frequency response, and phase response [smith1997scientist].
@@ 4.7.2 Poles and Zeros of the System Function
The poles and zeros of the system function $H(z)$ are essential for understanding the behavior of the system [oppenheim1997signals]. Zeros are the values of $z$ that make $H(z)=0$, and poles are the values of $z$ that make $H(z)$ undefined (where the denominator $D(z)=0$).
@@ 4.7.3 Stability of a System
The stability of an LTI system depends on the location of the poles of the system function $H(z)$ in the Z-plane. For a system to be stable, all poles must lie inside the unit circle ($|z|<1$) in the Z-plane [forouzan2016region]. If any pole lies outside or on the unit circle, the system is unstable.
@@ 4.8 Applications of Z-Transform
The Z-transform is a powerful mathematical tool used extensively in the analysis and design of digital systems [gonzalez2002digital]. It provides insights into the behavior of discrete-time systems in both the time and frequency domains. Here are a few key applications of the Z-transform in DSP [proakis1996digital].
@@ 4.8.1 Digital Filter Design
One of the primary applications of the Z-transform is in the design of digital filters [oppenheim2010dsp]. Filters are used to selectively enhance or suppress specific frequency components in a signal. By analyzing the Z-transform of the filter’s impulse response, we can design and analyze filters such as low-pass, high-pass, band-pass, and band-stop filters.
@@ 4.8.2 Control Systems Analysis
The Z-transform is widely used in the analysis of discrete-time control systems [mitra2002digital]. In control theory, the Z-transform helps analyze system stability, transient response, and steady-state behavior. By examining the system’s poles in the Z-plane, control engineers can assess whether the system will reach a stable equilibrium and how it will react to changes in input.
@@ 4.8.3 Time-Domain and Frequency-Domain Analysis
The Z-transform provides a bridge between the time domain and the frequency domain [oppenheim2010dsp]. By analyzing the Z-transform of a signal or system, engineers can understand how a system behaves over time and across different frequencies [proakis1996digital]. This is particularly useful in applications like audio processing, communications, and image compression, where signals are analyzed and processed in both domains [smith1997scientist].
@@ 5.1 Types of Digital Filters
Digital filters are essential components in digital signal processing, used to manipulate or extract information from discrete-time signals [oppenheim2010dsp]. They are algorithms or devices that operate on a digital input signal to produce a digital output signal. Digital filters are broadly categorized into two main types: Finite Impulse Response (FIR) filters and Infinite Impulse Response (IIR) filters. Each type has unique characteristics, advantages, and applications.
@@ 5.1.1 Finite Impulse Response (FIR) Filters
FIR filters have an impulse response that is of finite duration, meaning it settles to zero in finite time. The general form of an FIR filter is given by the convolution sum [proakis1996digital]:
@@
  • NN is the number of taps (filter order + 1).
@@
  • Always Stable: Since FIR filters do not have feedback elements, they are inherently stable [oppenheim2010dsp].
  • Linear Phase Response: FIR filters can be designed to have a linear phase response, which preserves the wave shape of signals within the passband [smith2007introduction].
  • Ease of Design: FIR filters are generally easier to design using standard methods [smith2007introduction].
@@ 5.1.2 Infinite Impulse Response (IIR) Filters
IIR filters have an impulse response that theoretically lasts forever, as they use feedback from output to input [oppenheim2010dsp]. The general form of an IIR filter is given by the difference equation:
@@
  • Potentially Unstable: Due to feedback, stability must be carefully considered [proakis1996digital].
  • Efficient: IIR filters can achieve a desired response with a lower filter order than FIR filters [smith2007introduction].
  • Non-linear Phase Response: IIR filters generally do not have linear phase characteristics [oppenheim2010dsp].
@@ 5.2 FIR Filter Design Techniques
Designing FIR filters involves determining the filter coefficients $b_{k}$ that meet specific frequency response characteristics. Two commonly used techniques for FIR filter design are the Window Method and the Frequency Sampling Method [proakis1996digital].
@@ 5.2.1 Window Method
The Window Method is a straightforward and widely used approach for designing FIR filters. The main idea is to truncate the infinite impulse response of an ideal filter to a finite length by applying a window function [smith2007introduction]. This process introduces ripples in the frequency response due to the Gibbs phenomenon, but by choosing appropriate window functions, we can control the trade-off between main-lobe width and side-lobe levels [proakis1996digital].
@@
The general steps for designing an FIR filter using the window method are as follows [smith2007introduction]:
@@
    Choose a Window Function: Select an appropriate window function $w[n]$ (e.g., Rectangular, Hamming, Hanning, Blackman, or Kaiser window) that meets the design specifications [proakis1996digital].
@@
Different window functions affect the frequency response differently [smith2007introduction]. Here are some commonly used windows:
@@
    Simplest window with poor frequency response (high side-lobes) [smith2007introduction].
@@
    Provides better side-lobe attenuation than the Rectangular window [smith2007introduction].
@@
    Similar to Hamming but with different side-lobe characteristics [smith2007introduction].
@@
    Offers better side-lobe attenuation at the expense of a wider main lobe [smith2007introduction].
@@
    Where $I_{0}$ is the zero-th order modified Bessel function, and $\beta$ is a parameter that controls the trade-off between main-lobe width and side-lobe level [proakis1996digital].
@@
We can use Python to compute the filter coefficients and visualize the frequency response [virtanen2020scipy].